# Revisiting Entropy in Reinforcement Learning for Large Reasoning Models

### v1

@@ 6.2 Training Data Diversity
From Table 2, we observe that entropy decreases as the amount of training data decreases. Notably, LLMs trained on subsets constructed via K-means clustering consistently exhibit lower entropy than those trained on randomly sampled subsets, except in the case of models trained on 5,031 samples. These results further demonstrate that training data diversity affects entropy dynamics, as entropy tends to decline when data diversity diminishes. Moreover, Table 2 reveals that, despite a substantial reduction in training data, LLMs trained on considerably smaller subsets can still achieve performance comparable to those trained on the full dataset. This finding indicates that data scale alone does not determine model performance, consistent with prior studies (DBLP:journals/corr/abs-2502-11886; DBLP:journals/corr/abs-2502-03387; DBLP:journals/corr/abs-2501-19393).
6.3 Clipping Threshold
A higher upper clipping threshold in GRPO helps alleviate entropy collapse, whereas a lower threshold tends to intensify it. A similar pattern holds for the lower clipping threshold: increasing it mitigates entropy collapse, while decreasing it amplifies the effect. Notably, LLMs can be trained stably even without clipping, achieving performance comparable to that obtained with clipping.
To comprehensively examine how clipping thresholds affect the entropy and performance of LLMs, we trained LLMs with RLVR under various lower and upper clipping thresholds beyond the default setting of $\varepsilon_{\text{low}}=\varepsilon_{\text{high}}=0.2$. Specifically, we conducted the following experiments:
  • Clip-Higher (DBLP:journals/corr/abs-2503-14476) raises the upper clipping bound in the GRPO objective to reduce the proportion of low-probability tokens being clipped. This relaxation allows these tokens to increase their likelihoods more freely, thereby enhancing exploration and mitigating entropy collapse in LLMs. Following DBLP:journals/corr/abs-2503-14476, we set $\varepsilon_{\text{high}}=0.28$.
  • Clip-Tighter, in contrast to Clip-Lower, decreases $\varepsilon_{\text{low}}$, thereby raising the clipping lower bound in the GRPO objective. As a result, low-probability tokens with negative advantages become more likely to be clipped, preventing their probabilities from being excessively suppressed. Hence, Clip-Tighter is expected to alleviate entropy collapse in LLMs. Specifically, while Clip-Lower increases $\varepsilon_{\text{low}}$ from its default value of 0.2 to 0.28, Clip-Tighter decreases $\varepsilon_{\text{low}}$ by 0.08, resulting in a final value of 0.12.
The entropy dynamics of LLMs trained with RLVR under various lower and upper clipping thresholds are illustrated in Figure 9. As shown, Clip-Higher effectively prevents entropy collapse and even leads to an increase in model entropy during training. In contrast, the other clipping variants (Clip-Lower, Clip-Tighter, and Clip-Free) induce varying degrees of entropy collapse. Among them, Clip-Lower results in the most pronounced collapse, whereas Clip-Tighter mitigates entropy collapse and maintains higher entropy compared to the default clipping configuration ($\varepsilon_{\text{low}}=\varepsilon_{\text{high}}=0.2$).
Figure 9: Evolution of the entropy of LLMs during RLVR training with various lower and upper clipping thresholds.
Surprisingly, Figure 9 shows that Clip-Free attains the highest entropy among all clipping variants (including the default clipping configuration), except for Clip-Higher. Moreover, as presented in Table 3, on in-domain test sets, Clip-Free yields the second-highest average Avg@64 and Pass@64 scores among all clipping variants. On out-of-domain test sets, Clip-Free similarly achieves the second-highest Avg@64 and the highest average Pass@64. These results suggest that the clipping operation in the GRPO objective, which is designed to prevent excessively large updates during training, can be removed without compromising the stability of RLVR training.
    + Prog-Adv-Reweight-1 31.72 / 56.67 15.21 / 46.67 78.79 / 96.00 64.84 / 95.00 33.67 / 65.81 4.92 / 31.99 33.07 / 60.91 44.85 / 72.03 19.00 / 46.45 0.32983
    + Prog-Adv-Reweight-2 32.34 / 66.67 17.45 / 43.33 75.75 / 95.40 66.17 / 95.00 33.51 / 60.29 6.93 / 33.46 31.63 / 66.43 45.05 / 72.14 19.28 / 49.94 0.05203
Table 3: Performance of Qwen2.5-Math-7B trained with GRPO and its entropy-regularized variants. “Ada-Ent-Reg” denotes Adaptive Entropy Regularization. “Average (ID)” and “Average (OOD)” indicate the mean performance across in-domain and out-of-domain benchmarks, respectively. Results are presented as A / B, representing Avg@64 and Pass@64, evaluated with temperature = 1.0, top-$p$ = 1.0, and a maximum generation length of 4096 tokens.
To effectively regulate entropy, we first examine which types of tokens that drive entropy changes during training. zhihu2025entropy and DBLP:journals/corr/abs-2505-22617 mathematically proved that the entropy change of an LLM between two consecutive training steps can be approximated by the product of the learning rate and the covariance between $\log\pi_{\theta}^{k}\left(\bm{y}_{t}|\bm{x},\bm{y}_{<t}\right)$ and $\pi_{\theta}^{k}\left(\bm{y}_{t}|\bm{x},\bm{y}_{<t}\right)\hat{A_{t}}$. This implies that tokens with both high probabilities and high advantages, or low probabilities and low advantages, primarily contribute to the entropy collapse. However, we posit that tokens with low probabilities and low advantages occur infrequently, as high-probability tokens are more likely to be sampled during decoding and thus dominate model-generated responses. Based on this intuition, we hypothesize that the entropy collapse of LLMs mainly stems from tokens with high advantages, and we further provide a theoretical derivation to support this hypothesis.
Let $z_{v}$ denote the logit of token $v$ produced by the LLM. When token $v$ is not sampled in the generation of token $\bm{y}_{t}$, the gradient of the GRPO objective with respect to $z_{v}$ can be approximated as:
$\frac{\partial\mathcal{J}\left(\theta\right)}{\partial z_{v}}$ $=-\mathbb{E}_{\bm{y}_{t}\sim\pi_{\theta}\left(\cdot|\bm{x},\bm{y}_{<t}\right)}\pi_{\theta}\left(v|\bm{x},\bm{y}_{<t}\right)\hat{A}_{t}$ (7)
$\approx-\pi_{\theta}\left(v|\bm{x},\bm{y}_{<t}\right)\hat{A}_{t}$
Similarly, when token $v$ is sampled in generating token $\bm{y}_{t}$, the gradient of the GRPO objective with respect to $z_{v}$ can be approximated as:
$\frac{\partial\mathcal{J}\left(\theta\right)}{\partial z_{v}}$ $=\mathbb{E}_{\bm{y}_{t}\sim\pi_{\theta}\left(\cdot|\bm{x},\bm{y}_{<t}\right)}\left(1-\pi_{\theta}\left(v|\bm{x},\bm{y}_{<t}\right)\right)\hat{A}_{t}$ (8)
The detailed derivations of the above approximations are provided in Appendix C. These approximations theoretically indicate that tokens with positive advantages drive the entropy collapse of LLMs during training. Specifically, since RLVR performs gradient ascent to maximize the optimization objective, a positive advantage implies that the update increases the probabilities of sampled tokens while decreasing those of unsampled ones. Given that high-probability tokens are more likely to be sampled, this process concentrates the probability mass and thus exacerbates entropy collapse. Conversely, when the advantages are negative, the update reduces the probabilities of sampled tokens and increases those of unsampled tokens. In this case, because high-probability tokens are again more likely to be sampled, the update counteracts the tendency toward over-concentration of probabilities, thereby alleviating entropy collapse.
To empirically validate this hypothesis, we conducted two comparative experiments: training LLMs solely on tokens with advantages $\geq 0$ and solely on tokens with advantages $\leq 0$, and compared these settings with the following baselines:
  • Adaptive Entropy Regularization dynamically adjusts the entropy regularization coefficient to keep model entropy above a predefined threshold $\delta$. We consider two settings for $\delta$: (1) $\delta=0.2$, following DBLP:journals/corr/abs-2505-22312, and (2) $\delta$ set to the entropy of the LLM on responses to 1,000 randomly sampled training prompts before training, which equals 0.3657.
  • Clip-Cov (DBLP:journals/corr/abs-2505-22617) mitigates entropy collapse by zeroing the gradients of a small subset of tokens with high covariance between log probabilities and advantages.
  • KL-Cov (DBLP:journals/corr/abs-2505-22617) adopts a similar approach to Clip-Cov but applies a KL penalty to high-covariance tokens.
  • Entropy-Adv (DBLP:journals/corr/abs-2506-14758) augments the original advantage with an entropy term to encourage exploratory reasoning tokens and enhance LLM performance.
As shown in Section 5, increasing the number of parameter updates per batch ($N_{\text{update}}$) intensifies entropy collapse. Accordingly, we primarily conduct experiments under the setting where $N_{\text{update}}$ is fixed at 4 to evaluate the effectiveness of different approaches in controlling the entropy of LLMs.
Figure 10 illustrates the evolution of entropy across different methods. As shown, Clip-Cov, KL-Cov, Adaptive Entropy Regularization, and Rand-Pos-Clip effectively alleviate entropy collapse, whereas Ent-Adv intensifies it. Furthermore, we observe that training LLMs exclusively on tokens with advantages $\geq 0$ leads to the most severe entropy collapse, while training on tokens with advantages $\leq 0$ drives entropy to a high level. These findings empirically support our hypothesis that entropy collapse primarily stems from tokens with high advantages, implying that adjusting the relative loss weights of tokens with advantages $\geq 0$ and $\leq 0$ can regulate model entropy.
Building on this insight, we propose Progressive Advantage Reweighting (Prog-Adv-Reweight), which controls entropy by dynamically adjusting the loss weights assigned to tokens with non-negative advantages. Specifically, we introduce a hyperparameter $\lambda$ that determines the loss weights of tokens with advantages $\geq 0$. We propose two variants of Prog-Adv-Reweight, distinguished by the way the $\lambda$ changes during training:
  • Prog-Adv-Reweight-1 partitions the RLVR training process into two stages, each comprising an equal number of training steps. In the first stage, $\lambda$ is fixed at 0, meaning the model is trained solely on tokens with non-positive advantages. In the second stage, $\lambda$ increases linearly from 0 to 1, progressively enhancing the contributions of positive-advantage tokens as training advances.
  • Prog-Adv-Reweight-2 increases $\lambda$ linearly across epochs, from 0 in the first epoch to 1 in the final epoch. If the total number of epochs is $E$ and the current epoch is $e$, then $\lambda$ is defined as $\lambda=(e-1)/(E-1)$.
As shown in Figure 10, both Prog-Adv-Reweight-1 and Prog-Adv-Reweight-2 induce an initial rise in entropy followed by a gradual decline, indicating that entropy decreases as the loss weights of tokens with positive advantages increase. This observation confirms that adjusting the relative loss weights of tokens with positive and negative advantages effectively regulates entropy.
Table 3 summarizes the performance of LLMs trained with RLVR under various entropy regularization approaches across multiple benchmarks. As shown in Table 3, although training exclusively on tokens with advantages $\leq 0$ effectively mitigates entropy collapse, it results in lower average Avg@64 scores on both in-domain and out-of-domain benchmarks, as well as reduced average Pass@64 scores on out-of-domain benchmarks. In contrast, by dynamically adjusting the loss weights of tokens with non-negative advantages during training, both Prog-Adv-Reweight-1 and Prog-Adv-Reweight-2 further improve LLM performance beyond the $\text{Adv}\leq 0$ setting and achieve results comparable to other entropy regularization methods. These findings demonstrate that Progressive Advantage Reweighting effectively mitigates entropy collapse while maintaining competitive performance. Moreover, despite its simplicity, Rand-Pos-Clip, which randomly sets the gradients of a small subset of tokens with positive advantages to zero, achieves an average Avg@64 score comparable to that of Clip-Cov on both in-domain and out-of-domain benchmarks, and a comparable average Pass@64 score on out-of-domain benchmarks.
8 Conclusion
In this paper, we comprehensively investigate the entropy dynamics of LLMs trained with RLVR. Through extensive empirical analyses, we examine how the entropy of LLMs correlates with response diversity, model calibration, and performance across multiple benchmarks. Furthermore, we identify three key factors that critically influence the entropy dynamics of LLMs: the number of off-policy updates, training data diversity, and the clipping threshold. Our theoretical and empirical analyses further reveal that entropy collapse primarily arises from tokens with positive advantages, and that entropy can be effectively regulated by adjusting the relative loss weights assigned to tokens with positive and negative advantages. Building on these insights, we propose Progressive Advantage Reweighting (Prog-Adv-Reweight), a simple yet effective approach that dynamically adjusts the loss weights of tokens with positive advantages during training, thereby mitigating entropy collapse while maintaining competitive performance compared with other entropy regularization methods across benchmarks.
Limitations
Due to computational constraints, our experiments were conducted solely on Qwen2.5-Math-7B. As LLMs vary in their pre-training data, training strategies, and architectures, some of our findings may not generalize directly to other models. Nevertheless, we hope that our analysis of entropy dynamics in RLVR training provides valuable empirical insights for the research community and motivates the development of more effective entropy regularization methods in future work.
Appendix A Experimental Setup
The diversity of responses generated by LLMs is evaluated using the N-gram Diversity (DBLP:conf/naacl/LiGBGD16) and SelfBLEU (DBLP:conf/sigir/ZhuLZGZWY18) metrics. Specifically, the N-gram Diversity metric quantifies the proportion of unique n-grams relative to the total number of n-grams in the generated responses. Let $U_{i}$ denote the number of unique n-grams and $C_{i}$ the total number of n-grams of order $i$. The metric is formally defined as
$\text{N-gram Diversity}=\prod_{i=1}^{N}\frac{U_{i}}{C_{i}}.$ (9)
In our experiments, we set $N=5$ when applying the N-gram Diversity metric to assess response diversity.
The SelfBLEU metric measures diversity from a complementary perspective. For each response, the BLEU score is computed by treating it as the hypothesis and all other responses as references. The resulting BLEU scores, averaged over 1- to 4-gram, are then averaged across all responses to obtain the final SelfBLEU score.
Appendix B How Does the Entropy of LLMs Trained with RLVR Correlate with Their Performance?
Figure 11: Evolution of entropy (solid lines) and negative SelfBLEU (markers) over training steps under different entropy regularization settings.
Appendix C How Can Entropy Be Effectively Regulated to Improve the Performance of LLMs?

### v2

@@ D.3 Training Data Diversity
Appendix E How Can Entropy Be Effectively Regulated to Improve the Performance of LLMs?
E.1 Theoretical Analysis
Figure 12: Evolution of entropy and training rewards under the Clip-Higher setting.
@@ E.1 Theoretical Analysis
E.2 Empirical Analysis
The baseline methods against which we compare are described as follows:
  • Adaptive Entropy Regularization dynamically adjusts the entropy regularization coefficient to keep model entropy above a predefined threshold $\delta$. We consider two settings for $\delta$: (1) $\delta=0.2$, following He et al. (2025), and (2) $\delta$ set to the entropy of the LLM on responses to 1,000 randomly sampled training prompts before training, which equals 0.3657.
  • Clip-Cov (Cui et al., 2025) mitigates entropy collapse by zeroing the gradients of a small subset of tokens with high covariance between log probabilities and advantages.
  • KL-Cov (Cui et al., 2025) adopts a similar approach to Clip-Cov but applies a KL penalty to high-covariance tokens.
  • Entropy-Adv (Cheng et al., 2025) augments the original advantage with an entropy term to encourage exploratory reasoning tokens and enhance LLM performance.
We conduct experiments with $N_{\text{update}}$ fixed at 4 to evaluate the effectiveness of different approaches in controlling the entropy of LLMs.
E.3 Analysis of the Covariance between Log-Probability and Advantage
Figure 13: Evolution of the negative entropy change and the covariance terms during GRPO training of Qwen2.5-Math-7B with $N_{\text{update}}=1$. The left figure reports the covariance between $\log\pi_{\theta}^{k}\left(\bm{y}_{t}|\bm{x},\bm{y}_{<t}\right)$ and $\pi_{\theta}^{k}\left(\bm{y}_{t}|\bm{x},\bm{y}_{<t}\right)\hat{A}_{t}$, while the right panel reports the covariance between $\log\pi_{\theta}^{k}\left(\bm{y}_{t}|\bm{x},\bm{y}_{<t}\right)$ and $\hat{A}_{t}$.
Figure 14: Evolution of the negative entropy change and the covariance terms during GRPO training of Qwen2.5-Math-7B with $N_{\text{update}}=4$. The left figure reports the covariance between $\log\pi_{\theta}^{k}\left(\bm{y}_{t}|\bm{x},\bm{y}_{<t}\right)$ and $\pi_{\theta}^{k}\left(\bm{y}_{t}|\bm{x},\bm{y}_{<t}\right)\hat{A}_{t}$, while the right panel reports the covariance between $\log\pi_{\theta}^{k}\left(\bm{y}_{t}|\bm{x},\bm{y}_{<t}\right)$ and $\hat{A}_{t}$.
Liu (2025) and Cui et al. (2025) provide a theoretical analysis showing that the change in entropy of an LLM between two consecutive training steps is governed by the covariance between token log-probabilities and advantages. Formally, let $\eta$ denote the learning rate at training step $k$, and let $\bm{y}_{t}$ be a token sampled from the policy $\pi_{\theta}^{k}\left(\bm{y}_{t}|\bm{x},\bm{y}_{<t}\right)$. Under the policy gradient, the entropy change of the policy $\pi_{\theta}$ from step $k+1$ to $k$ can be approximated as follows:
$\mathcal{H}\left(\pi_{\theta}^{k+1}\left(\bm{y_{t}}|\bm{x},\bm{y}_{<t}\right)\right)-\mathcal{H}\left(\pi_{\theta}^{k}\left(\bm{y}_{t}|\bm{x},\bm{y}_{<t}\right)\right)\approx$ (16)
$-\eta\cdot\text{Cov}\left(\log\pi_{\theta}^{k}\left(\bm{y}_{t}|\bm{x},\bm{y}_{<t}\right),\pi_{\theta}^{k}\left(\bm{y}_{t}|\bm{x},\bm{y}_{<t}\right)\hat{A}_{t}\right).$
Figure 15: Evolution of the negative entropy change and the covariance terms during GRPO training of Qwen2.5-Math-7B with $N_{\text{update}}=4$, where updates are performed exclusively on tokens with $\text{Adv}\leq 0$. The left figure reports the covariance between $\log\pi_{\theta}^{k}\left(\bm{y}_{t}|\bm{x},\bm{y}_{<t}\right)$ and $\pi_{\theta}^{k}\left(\bm{y}_{t}|\bm{x},\bm{y}_{<t}\right)\hat{A}_{t}$, while the right panel reports the covariance between $\log\pi_{\theta}^{k}\left(\bm{y}_{t}|\bm{x},\bm{y}_{<t}\right)$ and $\hat{A}_{t}$.
Figure 16: Evolution of the negative entropy change and the covariance terms during GRPO training of Qwen2.5-Math-7B with $N_{\text{update}}=4$, where updates are performed exclusively on tokens with $\text{Adv}\geq 0$. The left figure reports the covariance between $\log\pi_{\theta}^{k}\left(\bm{y}_{t}|\bm{x},\bm{y}_{<t}\right)$ and $\pi_{\theta}^{k}\left(\bm{y}_{t}|\bm{x},\bm{y}_{<t}\right)\hat{A}_{t}$, while the right panel reports the covariance between $\log\pi_{\theta}^{k}\left(\bm{y}_{t}|\bm{x},\bm{y}_{<t}\right)$ and $\hat{A}_{t}$.
Similarly, under the natural policy gradient, the entropy change between steps $k$ and $k+1$ can be expressed as follows:
$\mathcal{H}\left(\pi_{\theta}^{k+1}\left(\bm{y_{t}}|\bm{x},\bm{y}_{<t}\right)\right)-\mathcal{H}\left(\pi_{\theta}^{k}\left(\bm{y}_{t}|\bm{x},\bm{y}_{<t}\right)\right)\approx$ (17)
$-\eta\cdot\text{Cov}\left(\log\pi_{\theta}^{k}\left(\bm{y}_{t}|\bm{x},\bm{y}_{<t}\right),\hat{A}_{t}\right).$
To connect our empirical analysis with these theoretical derivations, we reran the experiments on Qwen2.5-Math-7B to record both $\text{Cov}(\log\pi_{\theta}^{k}(\bm{y}_{t}|\bm{x},\bm{y}_{<t}),\pi_{\theta}^{k}(\bm{y}_{t}|\bm{x},\bm{y}_{<t})\hat{A}_{t})$ and $\text{Cov}(\log\pi_{\theta}^{k}(\bm{y}_{t}|\bm{x},\bm{y}_{<t}),\hat{A}_{t})$ throughout training. Specifically, we considered the following settings:
  • GRPO with $N_{\text{update}}=1$;
  • GRPO with $N_{\text{update}}=4$;
  • GRPO with $N_{\text{update}}=4$ trained exclusively on tokens with $\text{Adv}\leq 0$;
  • GRPO with $N_{\text{update}}=4$ trained exclusively on tokens with $\text{Adv}\geq 0$;
  • GRPO with $N_{\text{update}}=4$ augmented with Pos-Adv-Reweight (Entropy-guided).
Figure 13 to 17 illustrates the evolution of the negative entropy change and the corresponding covariance terms under the above settings, as well as the Spearman’s rank correlation coefficients between the negative entropy change and each covariance term. As shown in these figures, both the negative entropy change and the covariance terms exhibit noticeable fluctuations during training, while the absolute values of the Spearman’s rank correlation coefficients remain generally small. This observation suggests that the relationship between entropy change and these covariance terms may be influenced by multiple factors, including the training data and the LLMs employed for GRPO training.
Figure 17: Evolution of the negative entropy change and the covariance terms during GRPO training augmented with Pos-Adv-Reweight (Entropy-guided) on Qwen2.5-Math-7B with $N_{\text{update}}=4$. The left figure reports the covariance between $\log\pi_{\theta}^{k}\left(\bm{y}_{t}|\bm{x},\bm{y}_{<t}\right)$ and $\pi_{\theta}^{k}\left(\bm{y}_{t}|\bm{x},\bm{y}_{<t}\right)\hat{A}_{t}$, while the right panel reports the covariance between $\log\pi_{\theta}^{k}\left(\bm{y}_{t}|\bm{x},\bm{y}_{<t}\right)$ and $\hat{A}_{t}$.
Appendix F Experiments on Llama-3.1-8B-Instruct
F.1 Experimental Setup
To assess whether our empirical findings generalize beyond Qwen2.5-Math-7B, we further trained Llama-3.1-8B-Instruct using GRPO. [^1] We initially intended to train Llama-3.1-8B with GRPO, as it is a pretrained model that has not undergone instruction tuning, similar to Qwen2.5-Math-7B. However, during preliminary experiments, we observed that Llama-3.1-8B frequently generated endlessly repetitive responses during training, which substantially slowed down the training process and degraded training stability. Consequently, we selected Llama-3.1-8B-Instruct, which has undergone instruction tuning, to ensure more efficient and stable training. The training configuration for Llama-3.1-8B-Instruct followed that of Qwen2.5-Math-7B, as described in Section 4. Specifically, Llama-3.1-8B-Instruct was trained on the DAPO-Math-17K dataset (Yu et al., 2025) with the AdamW optimizer, employing a cosine learning rate schedule and a peak learning rate of $1\times 10^{-6}$. During rollout, the batch size was set to 256, and 16 responses were generated per prompt using top-$p$ sampling with $p=1.0$, a temperature of 1.0, and a maximum generation length of 4096 tokens. For evaluation, we largely followed the setup used for Qwen2.5-Math-7B, with one exception: MATH500 was used as the validation set. This choice was motivated by the generally poor performance of GRPO-trained Llama-3.1-8B on AIME 2025, which makes AIME 2025 an unreliable validation benchmark in this setting.
Figure 18: Evolution of entropy (solid lines) and N-gram diversity (markers) across training steps under varying numbers of off-policy updates for Llama-3.1-8B-Instruct trained with GRPO.
Figure 19: Evolution of entropy (solid lines) and negative SelfBLEU scores (markers) across training steps under varying numbers of off-policy updates for Llama-3.1-8B-Instruct trained with GRPO.
F.2 Entropy and Response Diversity
Figure 18 and Figure 19 illustrate the entropy dynamics of Llama-3.1-8B-Instruct trained with GRPO under varying numbers of off-policy updates, together with the diversity of generated responses across training steps, measured by N-gram Diversity and SelfBLEU, respectively. As shown in Figure 18, increasing the number of off-policy updates leads to a reduction in model entropy during training, which is accompanied by a corresponding decline in the diversity of the generated responses. Notably, a strong positive correlation is observed between entropy and N-gram Diversity, with a Spearman’s rank correlation coefficient of 0.8937. This result indicates that the empirical relationship between entropy and response diversity identified for Qwen2.5-Math-7B in Section C.1 also holds for Llama-3.1-8B-Instruct.
F.3 Entropy Dynamics on Prompts
Figure 20 presents the ratio between the entropy of Llama-3.1-8B-Instruct measured at different training steps and its initial entropy prior to training, across different numbers of off-policy updates. As illustrated in Figure 20, the entropy measured on prompts drawn from the same domain as the training data decreases more rapidly than that measured on prompts from different domains. A similar trend is observed for Qwen2.5-Math-7B, indicating that the empirical findings reported in Section C.2 also hold for Llama-3.1-8B-Instruct.
F.4 Prompt Entropy and Accuracy
Figure 21 illustrates the relationship between accuracy and entropy for Llama-3.1-8B-Instruct across AIME 2024, AIME 2025, and MATH500. As shown in the figure, the entropy measured on prompts exhibits only a weak correlation with the accuracy of the corresponding responses. Quantitatively, the average Spearman’s rank correlation coefficient between prompt entropy and response accuracy across the three benchmarks is 0.2078. Furthermore, Figure 22 shows that the Spearman’s rank correlation coefficient remains consistently small throughout GRPO training. These results indicate that the empirical finding reported in Section C.3, namely that prompt entropy is only weakly correlated with model accuracy, also holds for Llama-3.1-8B-Instruct.
Figure 20: Ratio of the entropy of Llama-3.1-8B-Instruct at different training steps to its initial entropy prior to training, under varying numbers of off-policy updates.
Figure 21: Scatter plot of accuracy versus entropy for Llama-3.1-8B-Instruct evaluated on AIME 2024, AIME 2025, and MATH500.
Figure 22: Spearman’s rank correlation coefficients between entropy and accuracy for Llama-3.1-8B-Instruct, measured at checkpoints saved at different training steps. “Average” indicates the mean Spearman’s rank correlation coefficient computed over AIME 2024, AIME 2025, and MATH500.
F.5 Correlations Between Entropy and Model Performance
Table 4 reports the Spearman’s rank correlation coefficients between model entropy and the Avg@64 and Pass@64 performance metrics for Llama-3.1-8B-Instruct trained with GRPO and its variants. As shown in Table 4, entropy exhibits a strong negative correlation with Avg@64 on IF-Eval and a strong negative correlation with Pass@64 on LiveCodeBench. These results indicate that the empirical finding obtained with Qwen2.5-Math-7B in Section 5.2, namely that the relationship between LLM entropy and benchmark performance depends on both the task and the evaluation metric, also holds for Llama-3.1-8B-Instruct.
Avg@64 Pass@64
AIME 2024 -0.08571 0.39466
AIME 2025 -0.34786 -0.46291
MATH500 -0.02857 -0.31429
AMC 2023 0.02857 -0.26482
Minerva -0.48571 -0.11595
LiveCodeBench -0.54286 -0.88571
IF-Eval -0.82857 0.46382
Table 4: Spearman’s rank correlation coefficients between the entropy of LLMs and their performance across different benchmarks. Coefficients with the largest absolute values are highlighted in bold.
Figure 23: Distribution of log probabilities for correct and incorrect responses produced by Llama-3.1-8B-Instruct and its GRPO-trained variants. In each violin plot, the black line denotes the mean value.
F.6 Entropy Collapse and Miscalibration
Figure 23 presents the distributions of log probabilities, as computed by the corresponding LLMs, for correct and incorrect responses generated by Llama-3.1-8B-Instruct and its GRPO-trained variants. As illustrated in Figure 23, responses produced by Llama-3.1-8B-Instruct trained with GRPO and its variants are assigned higher probabilities than those produced by Llama-3.1-8B-Instruct. This observation suggests that GRPO training increases the model’s confidence in the responses it generates.
Moreover, consistent with the empirical findings for Qwen2.5-Math-7B, Llama-3.1-8B-Instruct assigns higher average probabilities to correct responses than to incorrect ones, and this pattern also holds for its GRPO-trained variants. However, after GRPO training, the gap in average probability between correct and incorrect responses becomes smaller, indicating that Llama-3.1-8B-Instruct becomes increasingly miscalibrated.
Llama-3.1-8B-Instruct 2.34 / 33.33 0.26 / 13.33 38.50 / 84.60 15.98 / 87.50 16.06 / 53.31 8.36 / 30.88 74.42 / 96.16 14.63 / 54.42 41.39 / 63.52 N/A
+ GRPO ($N_{\text{update}}=1$) 8.85 / 36.67 0.57 / 16.67 55.08 / 87.00 39.65 / 72.50 25.99 / 55.51 10.54 / 29.41 77.90 / 93.88 26.03 / 53.67 44.22 / 61.65 0.13570
+ GRPO ($N_{\text{update}}=2$) 7.29 / 43.33 0.47 / 13.33 52.33 / 90.60 25.55 / 80.00 25.61 / 59.93 9.61 / 31.62 78.06 / 95.08 22.25 / 57.44 43.84 / 63.35 0.04994
+ GRPO ($N_{\text{update}}=4$) 5.52 / 30.00 0.47 / 16.67 49.82 / 91.00 26.64 / 87.50 24.33 / 61.76 9.77 / 30.51 77.62 / 94.84 21.35 / 57.39 43.69 / 62.68 0.03781
      + $\text{Adv}\leq 0$ 3.13 / 40.00 0.52 / 16.67 43.17 / 89.80 18.52 / 80.00 20.19 / 58.46 8.46 / 29.04 76.00 / 95.80 17.10 / 56.98 42.23 / 62.42 2.18019
      + $\text{Adv}\geq 0$ 5.99 / 36.67 0.73 / 26.67 49.40 / 89.60 23.63 / 85.00 25.14 / 55.88 10.85 / 31.99 79.43 / 94.48 20.98 / 58.76 45.14 / 63.23 0.02846
      + Pos-Adv-Reweight (Entropy-guided) 7.03 / 36.67 0.31 / 13.33 50.16 / 89.40 26.88 / 87.50 23.78 / 58.46 9.90 / 30.15 77.27 / 94.84 21.63 / 57.07 43.59 / 62.50 0.19455
Table 5: Performance of Llama-3.1-8B-Instruct trained with GRPO and its variants.
Notably, when Llama-3.1-8B-Instruct is trained exclusively on tokens with no-positive advantages, which are expected to reduce the probabilities of responses generated during the rollout stage, the average probabilities of both correct and incorrect responses nonetheless increase. In this setting, the difference in average probability between correct and incorrect responses is larger than that observed under GRPO training, indicating that miscalibration is alleviated relative to GRPO. In contrast, when Llama-3.1-8B-Instruct is trained exclusively on tokens with non-negative advantages, the difference in average probability between correct and incorrect responses becomes even smaller than that observed under GRPO training, thereby further exacerbating miscalibration.
Taken together, these results demonstrate that the empirical finding reported in Section 5.3, namely that training LLMs with GRPO can induce miscalibration, also holds for Llama-3.1-8B-Instruct.
F.7 Effect of Off-Policy Updates
Figure 24: Evolution of entropy and reward on the training set during GRPO training of Llama-3.1-8B-Instruct.
Figure 24 illustrates the entropy dynamics and the evolution of training reward when Llama-3.1-8B-Instruct is trained with GRPO under different numbers of off-policy updates. As shown in Figure 24, increasing the number of off-policy updates from 1 to 2 results in a more rapid decrease in entropy, and further increasing the number to 4 accelerates the entropy reduction even further. In addition, a larger number of off-policy updates leads to higher rewards on the training set for Llama-3.1-8B-Instruct trained with GRPO. In contrast, the test performance measured by the Avg@64 score deteriorates as the number of off-policy updates increases. These results indicate that the empirical findings discussed in Section 6.2 also apply to Llama-3.1-8B-Instruct.
F.8 Experimental Results of Pos-Adv-Reweight
To evaluate the effectiveness of Pos-Adv-Reweight on Llama-3.1-8B-Instruct, we train the model using GRPO with four off-policy updates, incorporating Pos-Adv-Reweight (Entropy-guided) under the same experimental setup as Qwen2.5-Math-7B. The entropy dynamics of Llama-3.1-8B-Instruct trained with GRPO and its variants are shown in Figure 25, and the corresponding performance on multiple benchmarks is reported in Table 5.
As illustrated in Figure 25, Llama-3.1-8B-Instruct trained with GRPO, as well as the variant trained exclusively on tokens with non-negative advantages, both suffer from entropy collapse. In contrast, training exclusively on tokens with non-positive advantages results in pronounced entropy fluctuations. Conversely, training with Pos-Adv-Reweight (Entropy-guided) maintains the model entropy at a stable level of approximately 0.2 during training. Furthermore, the results in Table 5 indicate that restricting training to either non-negative-advantage tokens or non-positive-advantage tokens leads to inferior average Avg@64 scores on in-domain benchmarks compared with GRPO. By contrast, Llama-3.1-8B-Instruct trained with Pos-Adv-Reweight (Entropy-guided) outperforms GRPO in terms of average Avg@64 scores. These results further demonstrate that Pos-Adv-Reweight, which adaptively adjusts the loss weights of tokens with positive advantages, effectively controls the entropy of LLMs while improving their overall performance.
Figure 25: Evolution of entropy for Llama-3.1-8B-Instruct trained with GRPO and its variants.